When two given functions, 𝑓(π‘₯) = π‘₯ 2 + 3 π‘Žπ‘›π‘‘ 𝑔(π‘₯) = 2π‘₯ βˆ’ 5, are divided with each other, a new third function is obtained. Find (𝑓/𝑔)(π‘₯) and identify its domain


1
Expert's answer
2021-09-27T23:28:13-0400

When two given functions, "\ud835\udc53(\ud835\udc65) = \ud835\udc65^2 + 3" and "\ud835\udc54(\ud835\udc65) = 2\ud835\udc65 \u2212 5", are divided with each other, a new third function is obtained: "(𝑓/𝑔)(𝑥)=\frac{x^2+3}{2x-5}." Let us identify its domain. Since the denominator of "\frac{x^2+3}{2x-5}" is equal to zero in the point "x=2.5," we conclude that the domain of the function "(\ud835\udc53\/\ud835\udc54)(\ud835\udc65)" is "(-\infty,2.5)\cup(2.5,+\infty)."